OpenStudy (anonymous):

car accidents from 2000 to 2010 can be modeled by the quadratic C=-45x^2+1350x+16,695, where x=1 corresponds to 2000, x=2 corresponds to 2001, and so on until x=11 corresponds to 2010 find the number of cumulative car accidents in 2005

OpenStudy (anonymous):

In 2005, x = 6. Plug this into the quadratic equation: C = -45(6)^2 + 1,350(6) + 16,695 C = -1,620 + 8,100 + 16,695 So in 2005, C = 23,175